OpenStudy (anonymous):

Find a function whose graph is a parabola with vertex (1,-2) and that passes through the point (4,16).

OpenStudy (anonymous):

Use the vertex formula for a parabola: y = a(x-h)^2 + k , where (h,k) is the vertex y = a(x-1)^2 -2 , now plug in x=4 and y=16 into this equation and solve for a. Then, you are done!

OpenStudy (anonymous):

I tried this method out but it doesnt concord with the answer given in the answer key.

OpenStudy (anonymous):

16 = a(4 - 1)² + (- 2) 16 = a(3)² - 2 16 = a(9) - 2 16 = 9a - 2 9a = 16 + 2 9a = 18 a = 18 / 9 a = 2 you should get this...

OpenStudy (anonymous):

Then from there how would you obtain an equation?

OpenStudy (anonymous):

so the function would be y=2(x-1)^2-2?

OpenStudy (anonymous):

Thats the part i didnt get but here the answer key says 2x^2 -4x

OpenStudy (anonymous):

maybe they just simplifyed the function?

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

im pretty sure thats what they did!

OpenStudy (anonymous):

Yea i just realized that now thanks anyways stupidity at its finest haah (from my part)

OpenStudy (anonymous):

its fine! good job tho! :)